Five years ago Jacob was one-sixth the age of his brother. Three years time his age doubled will match his brother's age. How old is Jacob now?

Answer by htmentor(1343)

Let J = Jacob's age now
Let B = his brother's age now
5 years ago Jacob was 1/6 his brother's age:
J-5 = (B-5)/6
In 3 years, his age doubled will equal his brother's age:
2*(J+3) = B+3
We have two equations and two unknowns
Use one of the equations to solve for B in terms of J:
2J + 6 = B + 3
B = 2J + 3
J-5 = ((2J+3)-5)/6
J-5 = (2J-2)/6
6J-30 = 2J-2
This gives J = 7
So Jacob's age is 7
Check:
His brother's age = B = 2*7 + 3 = 17
5 years ago Jacob was 2, his brother was 12 (=6 times his age)
3 years from now Jacob will be 10
2*10 = 20 = 17 + 3